OpenStudy (anonymous):

The interest that Paula earns on her investment is given by the equation l = 36t where t is the time in years that her money is invested. She makes a table to determine how much money she would earn for different values of t. If one of the values for t she selects is 5, what would her corresponding entry be for l? 365 41 7.2 180

OpenStudy (nirmalnema):

36*5=i 180=i

OpenStudy (anonymous):

thanks @nirmalnema

OpenStudy (nirmalnema):

welcome..!!

OpenStudy (anonymous):

Calib received $50 for his birthday that he put in his piggy bank. Each week, he puts 2 more dollars in his bank so that the amount of money in dollars D in his bank is given by D = 2w + 50 where w is the number of weeks. His sister Lara makes a table of reasonable values for w to determine how much money Calib will eventually have. If one of the values for w Lara selects is 14, what would be her corresponding entry for D? (Points : 1) 264 66 78 57

OpenStudy (nirmalnema):

2*14+50 78=d

OpenStudy (anonymous):

The interest that Cindy earns on her investment is given by the equation l = 42t where t is the time in years that her money is invested. She makes a table to determine how much money she would earn for different values of t. If one of the values for t she selects is 12, what would her corresponding entry be for l? (Points : 1) 54 504 4212 3.5 thanks by the way

OpenStudy (nirmalnema):

12*42=504

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Jose received $30 for his birthday that he put in his piggy bank. Each week, he puts 4 more dollars in his bank so that the amount of money in dollars D in his bank is given by D = 4w + 30 where w is the number of weeks. His sister Marisol makes a table of reasonable values for w to determine how much money Andre will eventually have. If one of the values for w Marisol selects is 36, what would be her corresponding entry for D? (Points : 1) 466 174 70 39

OpenStudy (anonymous):

4(16)+30=64+30=94